A hire tool firm finds that their net return from hiring tools is decreasing by 10% per annum. If their net gain on a certain tool this year is £400, find the possible total of all future profits from this tool (assuming the tool lasts forever).
Answer by jorel1380(3719)

If the net return from hiring tools is 10%, then the value of hiring tools is 90% of what it was the year before. If the sum of an infinite geometric series can be given by the equation: S = a / (1 - x), where S is the total, a is the amount started with, and x is the residual value of the item in question, we get:
S=400/(1-0.9)
S=400/0.1=£4000 as the total net gain from hiring the tools after infinity years. ☺☺☺☺